OpenStudy (anonymous):

Three hundred centimeters is the same as 3 kilometers 3 meters 3 millimeters 3 decimeters

OpenStudy (frostbite):

The prefix, centi, is always \(10^{-3}\). In other words: \(1 cm = 10^{-3} m\) or \(1 = 10^{-3} \frac{m}{cm}\) 300 can be written as \(3 \times 10^{3}\). So lets multiply 300 with 1 as we got above: \(\Large 10^{3} cm \times 10^{-3} \frac{m}{cm} = 10^{3-3} \frac{m \times cm}{cm}=3m \) See the idea?

OpenStudy (anonymous):

thanx

OpenStudy (frostbite):

Sorry I write wrong all the time. \[\large 3 \times 10^{3} cm \times 10^{-3} \frac{m}{cm} = 3 \times 10^{3-3} \frac{m \times cm}{cm}=3m\] Remember that \(10^{0}=1\)

OpenStudy (anonymous):

ok thanx

OpenStudy (frostbite):

Kein Problem.